OpenStudy (anonymous):

Create your own fraction with like bases, coefficients, and show its simplification

OpenStudy (anonymous):

You could do something like: \[\frac{ 9m^5 }{ 3m^3 }\] and then simplify the 9 and the 3 to 3 and 1 respectively and you could do that rule I told you before and simplify the m^5 and the m^3 to m^2 and 1 \[\frac{ 3m^2 }{ 1m^0 }=3m^2\]